OpenStudy (akonkel):

Morgan has two sisters. Together, their ages are three consecutive integers and add up to 48. What are the ages of Morgan and her sisters?

OpenStudy (anonymous):

guess !!

OpenStudy (akonkel):

i think it would be 14,16,&18 is this right?

OpenStudy (anonymous):

those numbers are not consecutive

OpenStudy (akonkel):

what does consecutive mean?

OpenStudy (anonymous):

they are consecutive EVEN numbers

OpenStudy (anonymous):

like \(3,4,5\) or \(10,11,12\) one after another

OpenStudy (akonkel):

oh I know what even nums are :) ok then hold a sec i think I got this

OpenStudy (akonkel):

15, 16, 17 right? @Nnesha

OpenStudy (anonymous):

\[15+16+17=48\]?

OpenStudy (anonymous):

yes of course you win

Nnesha (nnesha):

@akonel ^^^ its better to focus on one post :)

Nnesha (nnesha):

same thing here

OpenStudy (akonkel):

wait wat??? so is that correct they equal 48

Nnesha (nnesha):

you can add that at least trust on calculator

Nnesha (nnesha):

them**

OpenStudy (calculusxy):

x +(x+1) + (x+2) = 48 That way you can solve for x and then add either one or two, if necessary, to figure out the ages.

OpenStudy (akonkel):

Solve the inequality 3 - 2a > 7 (Points : 2) a < -2 a < 2 a > -2 a > 2 a = 2

OpenStudy (calculusxy):

This is pretty much like this: \[3 - 2a = 7\]

OpenStudy (calculusxy):

So once you figure out the value of "a" then just add on the sign ( > ) with the number, after the isolation.

Nnesha (nnesha):

well i will suggest u to keep this sign >

OpenStudy (anonymous):

hmmm not quite

Nnesha (nnesha):

bec when you divide or multiply both side by negative number you must haveto flip sign

OpenStudy (akonkel):

oh ok :)

Nnesha (nnesha):

yeah :)

OpenStudy (akonkel):

so would A=7